2 balls of masses 10g and 25g are thrown HORIZONTALLY from same height with velocities V and v ...............if they fall in time intervals T and t..................and V:v= 1:3....................find ratio of T:t PLS EXPLAIN!!!!!!!!!!

2 balls of  masses 10g and 25g are thrown HORIZONTALLY from same height with velocities V and v ...............if they fall in time intervals T and t..................and V:v= 1:3....................find ratio of T:t


PLS EXPLAIN!!!!!!!!!!

Grade:9

1 Answers

KARAN ACHARYA
12 years ago

T/t =1

reason: height is same. both have zero initial velocity in vertical direction.

            s=ut +1/2at.t  ,  s=uT+1/2atT.T

            substitute u=0

            => T=t